OpenStudy (anonymous):

Mikhail has a savings portfolio containing $12,000, of which $3,000 is invested in CDs. His friend, Ivan, has $2,000 invested in CDs, and this investment is proportional to Mikhail's. How much money is in Ivan's savings portfolio?

OpenStudy (anonymous):

3,000 / 12,000 = 1/4 or .25 2,000x = 1/4 x = 8,000

OpenStudy (alexanderphantomhive):

haha what she said >u<

OpenStudy (ddos_dragon):

Make sure to medal and close your question. Have a nice day!

OpenStudy (ddos_dragon):

By the way. You posted your question in the English section. There are plenty of good noggins to pick over in the math area. :3